OpenStudy (anonymous):

PLEASE HELP *MEDAL AND FAN* 1. Tom and his best friend Edward are seeking to join a gym. Tom saw on television that Platinum Gym was having a grand opening in a few months. Tom is considering a membership at Platinum Gym. Platinum Gym charges a one-time registration fee of $60, and membership is $20 a month. Tom’s best friend Edward says he plans to join Super Fit because Super Fit only charges $10 a month but has a one-time registration fee of $150.

OpenStudy (anonymous):

(a) Write an expression to represent the membership cost of Platinum Gym for x months. (b) Write an expression to represent the membership cost of Super Fit for x months. (c) If Tom decides to join Platinum Gym and Edward joins Super Fit, who would pay less during the first month of membership? Who would pay less after a year of membership? Show your work. (d) Is there a membership period for which Tom and Edward would pay the same amount? If so, how long would this membership be and how much would it cost? Show your work. (e) Which gym membership is a better deal? Explain why. Answer:

OpenStudy (anonymous):

@gerryliyana

OpenStudy (anonymous):

hold on, wait a minute :)

OpenStudy (anonymous):

Take your time!

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

a) y = 20x + 60 it means, for example, for 4 months he must pay y = 20*(4) + 60 y = 80 + 60 y = $140

OpenStudy (anonymous):

so an expression to represent the membership cost of Platinum Gym for x months is y = 20x + 60

OpenStudy (anonymous):

get it @Kyliejones ??

OpenStudy (anonymous):

Thanks so much!!

OpenStudy (anonymous):

did you get for b?

OpenStudy (anonymous):

No not yet. :(

OpenStudy (anonymous):

b) y = 10x + 150

OpenStudy (anonymous):

Thanks alot i wish i can hug you right now!

OpenStudy (anonymous):

aww hug me then .. :*

OpenStudy (anonymous):

-hugs you - There you go :) .

OpenStudy (anonymous):

how about c ?

OpenStudy (anonymous):

still havent figured out c or D or e

OpenStudy (anonymous):

are you still here?

OpenStudy (anonymous):

If Tom decides to join Platinum Gym and Edward joins Super Fit, who would pay less during the first month of membership? Answer: Tom, because the one-time registration fee of Platinum Gym ($60) is less than the one-time registration fee of Super Fit ( $150).

OpenStudy (anonymous):

i miss you @Kyliejones

OpenStudy (anonymous):

Back!

OpenStudy (anonymous):

Ill brb i gtg eat!

OpenStudy (anonymous):

Who would pay less after a year of membership? Answer: - Tom decides to join Platinum Gym So, after a year (12 months) of membership, Tom must pay y = 20x + 60 y = 20*(12) + 60 y = 240 + 60 y = $300 - Edward joins Super Fit So, after a year (12 months) of membership, Edward must pay y = 10x + 150 y = 10*(12) + 150 y = 120 + 150 y = $270

OpenStudy (anonymous):

get it baby?

OpenStudy (anonymous):

@Kyliejones

OpenStudy (anonymous):

I got it but im not your baby.

OpenStudy (anonymous):

ok., sorry

OpenStudy (anonymous):

its fine :)